VIBRATIONAL DYNAMICS AROUND THE CONICAL INTERSECTION RESULTING FROM THE Ã → X̃ LASER INDUCED FLUORESCENCE OF THE METHOXY (CH_{3}O) RADICAL

JAYASHREE NAGESH and \underline{EDWIN L. SIBERT III}, Department of Chemistry and Theoretical Chemistry Institute, University of Wisconsin-Madison, WI 53706.

The results of a theoretical calculation of the spectra associated with the laser induced fluorescence Ã^{2}A_{1} → X̃^{2}E of both the methoxy molecule and CH_{2}DO are presented and discussed. The form of the vibronic dipole moment is determined by symmetry and the corresponding dipole expansion coefficients are calculated using ab initio methods. The calculated spectra include states up to 3000 cm^{−1} above the zero point energy. We describe how the various features of the spectrum are related to coordinate dependent terms in the dipole expansion as well as the spin-orbit couplings, Jahn-Teller couplings, and vibrational anharmonicities.